if an angle is 30 more than one half of its complement, find the measure of the angle
Answer by mananth(12270)

let one angle be x
the compliment will be 90-x
half the compliment = 1/2 * (90-x)
x = 1/2*(90-x)+30
x =1/2 *90 - 1/2 *x +30
x =45 -1/2 *x +30
3x/2 = 75
3x = 150
x=50